OpenStudy (anonymous):

2^x+6=4 with steps please :)

OpenStudy (anonymous):

\[2^{x}+6= 4\]

OpenStudy (anonymous):

Is this your equation

OpenStudy (anonymous):

2^(x+6)

OpenStudy (anonymous):

Do you know how to write this as a logarithmic equation?

OpenStudy (anonymous):

YOu there?

OpenStudy (anonymous):

Yeah sorry my wifi is crappy. And no, I dont.

OpenStudy (anonymous):

\[\log_{a} b = c ....is. the. same. as.....a ^{c}=b..\]

OpenStudy (anonymous):

So what wold this look like as a logarithmic equation

OpenStudy (anonymous):

How do you know what a,b, and c are?

OpenStudy (anonymous):

Would it be x+6^4=2?

OpenStudy (anonymous):

A= 2 C= (x+6) B= 4 Sorry was disconnected. So what will the logarithmic equation look like?

OpenStudy (anonymous):

Log^2 4=x+6

OpenStudy (anonymous):

Correct now solve for\[\log_{2} 4\]

OpenStudy (anonymous):

Okay, sorry but how do you do that? I'm more of a science girl. Math isn't exactly my forte sadly. :(

OpenStudy (anonymous):

Do you have a calculator?

OpenStudy (anonymous):

Only on my phone.

OpenStudy (anonymous):

To determine \[\log_{2} 4\] you would dive log 4 by log 2 which equals 2.

OpenStudy (anonymous):

So , 2=x+6. Solve for x?

OpenStudy (anonymous):

You got it Science girl ;-}

OpenStudy (anonymous):

Thank you sooooo much! You just helped me understand this whole concept. ;D